Excrss pressure in advance drop of water of radius 2mm is 70  the pressure inside a drop of 4mm is
 
 

Grade:12

2 Answers

rushil
4 years ago
The excess pressure p is given by
2T, where T is surface tension and R, is the radius of the bubble
Given T=70×10-3 Nm-1
R=1mm=10-3 m
p=2T/R=2 x 70×10-3/10-3=140Nm-2
10
Khimraj
4 years ago
T=70×10-3 Nm-1
R=10-3 m
p=2T/R=2 x 70×10-3/10-3=140Nm-2
…............................................................